OpenStudy (anonymous):

Given : Area = 18720 D= 194 P= ? RECTANGLE .. HELP !

OpenStudy (anonymous):

@Loopki40

OpenStudy (anonymous):



OpenStudy (anonymous):

Help please

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

Ill give medal and ill fan those who help.

OpenStudy (anonymous):

I'll help you. :)

OpenStudy (anonymous):

Here's how I see, it: there are two triangles there.

OpenStudy (anonymous):

That means you can use 1/2bh = a

OpenStudy (anonymous):

So use that to find the lengths of each side.

OpenStudy (anonymous):

I am finding the a and b sir... I have the area of the rectangle but i am finding the perimeter of it.

OpenStudy (anonymous):

and the diagonal is 194

OpenStudy (anonymous):

OHHH lol

OpenStudy (anonymous):

The Diagonal is the square root of (width squared + height squared): d = sqrt(w^2 + h^2)

OpenStudy (anonymous):

so i have to find the a and b in order to compute the perimeter.

OpenStudy (anonymous):

Yes, but you'll have two equations that a and b will have to fit. d = sqrt(a^2 + b^2) and 2(a + b)

OpenStudy (anonymous):

Are you given any answer choices?

OpenStudy (anonymous):

nope i have to find it. our instructor just reverse the question he gave us . Its finding the area.

OpenStudy (anonymous):

so a = b * h

OpenStudy (anonymous):

you're given d, which is sqrt(a^2 + b^2)

OpenStudy (anonymous):

and you need to solve for p, which is 2(a + b)

OpenStudy (anonymous):

Gah, I can't see the connection. Do you?

OpenStudy (anonymous):

I think there is something wrong with the problem

OpenStudy (anonymous):

So do I.

OpenStudy (anonymous):

There isn't a calculator online that I can find that can use the variables given and find a and b.

OpenStudy (anonymous):

I think We have to derive a new formula but . I think there's no formula that we can derive because it is perimeter and the formula for the perimeter of a rectangle is PERIMETER = 2L+2W and the are is A=LXW

OpenStudy (triciaal):

Area of rectangle,A= a*b = 18720 perimeter, P = 2a + 2b a^2 + b^2 = 194^2 from the area we can substitute for a or b use this expression in the eqn with D to find b now that you have a and b calculate P

OpenStudy (anonymous):

No our teacher just reverse the question there is nothing wrong with the area

OpenStudy (anonymous):

What? Can you elaborate it please

OpenStudy (triciaal):

sorry I did not have D^2 which would be 194^2

OpenStudy (triciaal):

are you asking to go through the steps above?

OpenStudy (anonymous):

the perimeter is 548 but how can i get the a and b . I use the pythagorean theorem then i dont know what is next

OpenStudy (anonymous):

Steps please

OpenStudy (triciaal):

give me a minute i think i found a "short-cut"

OpenStudy (triciaal):

is the question still to find the perimeter or is the question changed and the perimeter is now 548 given

OpenStudy (triciaal):

are you there? I want to answer the right question?

OpenStudy (anonymous):

@triciaal hey the perimeter is 548

OpenStudy (anonymous):

Help please anyone

OpenStudy (anonymous):

anyone help please

OpenStudy (anonymous):

@imer

OpenStudy (ikram002p):

Given : Area = 18720 D= 194 b= ? since its regtangle then rectangle area=a*b  Hence :- ab=18720 and a^2+b^2=194^2